OpenStudy (babyx3boo):

(5x+1/5)^2 i got 25x^2+10 2/5x+1/25 and it keep saying it's wrong

alones (alones):

5x+15)2 =(5x+1/5)(5x+1/5) =(5x)(5x)+(5x)(1/5)+(1/5)(5x)+(1/5)(1/5) =25x2+x+x+1/25 =25x2+2x+1/25 (Decimal: 25x^{2}+2x+0.04)

OpenStudy (anitas.):

^ Alone is correct!!

OpenStudy (anitas.):

\[25x^2+2x+\frac{ 1 }{ 25 }\] is ur answer. Alone explained the steps!!

alones (alones):

YASS!<3

OpenStudy (babyx3boo):

thank you :)